# Towards interactive evaluations for interaction harms in human-AI systems

### v6

@@ 4 Towards better evaluations of interaction harms
These measurement targets can be assessed through both self-reported measures — including validated scales for attitudes, perceptions, and affective states — and behavioral measures that track concrete actions like response time, task accuracy, and engagement patterns (Damacharla et al., 2018; Coronado et al., 2022; Gordon et al., 2021). In Table 2, we present a non-exhaustive set of these metrics for assessing both human impact and human-AI performance.
@@
Psychological impact User affective states, or perceptions and attitudes toward the model or the interaction Usability metrics, user satisfaction surveys, psychometrically validated measures for a specific construct or harm (e.g., Decision Regret Scale (Brehaut et al., 2003))
@@
Principle 3: structure human participation to balance validity and practicality
@@
Chat logs from real-world interactions provide unparalleled ecological validity and scale. These naturally occurring conversations capture authentic user behaviors and emergent patterns that would be difficult to replicate in controlled settings (Tamkin et al., 2024). The massive volume of existing logs also enables analysis of rare events and edge cases that controlled studies might miss. Recent work has paired analyses of millions of chat logs of user conversations with ChatGPT with longitudinal surveys of those users to examine emotional well-being, showing a path forward for using chat logs while also assessing real user impact (Phang et al., 2025). However, these observational data is often proprietary, difficult to access, raises serious privacy concerns, and may limit our ability to test specific hypotheses or establish causal relationships (Reuel et al., 2024; Zhao et al., 2024; Ouyang et al., 2023).
@@
Existing research on user simulations has utilized LLMs to simulate believable human behaviors using a range of architectures from demographic-based models to simulations informed by qualitative interviews (Park et al., 2024, 2023). Recent work has used such simulations to evaluate safety risks, utilizing human experiments to validate the evaluations (Ibrahim et al., 2025; Zhou et al., 2024). Such simulations can be particularly valuable for exploring scenarios that would be ethically challenging to test with real users. Unlike real human participants, who naturally shape conversations in diverse ways, simulations also allow for controlled trajectories, making it easier to analyze how interactions evolve across different scenarios. However, their limitations in believability, accuracy, and diversity require further study, as recent work has shown that they misportray and flatten representations of marginalized identity groups (Wang et al., 2025; Agnew et al., 2024).
@@ 5 Open challenges and ways forward for interactive evaluations
In this paper, we motivate the need to investigate human-AI interaction dynamics that current evaluations miss. However, implementing such evaluations at scale presents several concrete challenges. These range from ethical questions about studying potentially harmful interactions, to practical needs for research infrastructure, to methodological questions about producing actionable insights for stakeholders. While our design principles outline key considerations for scenarios, measurements, and participation strategies, advancing these methods requires addressing key open questions. Here, we identify specific challenges and opportunities where concentrated research effort could significantly advance our ability to evaluate increasingly interactive AI systems.
@@
How can we ethically work with human participants on studying harms? When are user simulations appropriate replacements?
@@
Studying interaction harms poses inherent ethical challenges: we need to understand potentially harmful dynamics without exposing participants to those same harms. This requires careful consideration of when and how to involve human participants. For participant protection, AI researchers should adopt existing practices from fields experienced in providing adequate participant training, thoughtful debriefing, and active feedback collection to mitigate experimental harms (McKee, 2024). Simulated human interactions offer a promising direction for studying scenarios where minimizing harm is challenging, such as young people’s interactions with AI systems or emotional attachment to AI systems. Recent work suggests these simulations can believably represent human behavior (Park et al., 2023, 2024), but more work is needed to understand how to best capture diverse user behaviors and develop ethical guidelines for deploying user simulations in safety evaluations (Agnew et al., 2024; Anthis et al., 2025).
How can we improve researcher access to data for understanding interaction harms?
@@
Current public datasets of human-AI conversations, while valuable, fall short of meeting researchers’ needs. Large collections of chat logs like WildChat and ShareGPT offer millions of conversations but provide only a narrow window into AI usage, capturing self-selected interactions that may not represent typical user behaviors or patterns (Zhao et al., 2024; Ouyang et al., 2023). These datasets can be biased in terms of user demographics, use cases, and interaction styles, potentially leading to incomplete or incorrect conclusions about how people typically interact with AI systems. AI developers, such as large labs which provide commercial chatbot services, have access to the most comprehensive interaction data (Tamkin et al., 2024). While these labs have incentives to protect user privacy and proprietary information, developing privacy-preserving data sharing frameworks is necessary to enable researchers to access representative interaction data while maintaining user trust and consent.
What infrastructure do we need to facilitate interactive evaluations?
@@
Many interactive evaluations share common stages: recruiting participants, collecting interaction data, and administering surveys and other tasks. Developing accessible protocols, guides, and standardized test suites — similar to those available for non-interactive evaluations — can support and facilitate an increase in interactive evaluations (UK AI Safety Institute, 2024; METR, 2023; Collins et al., 2023). Some existing platforms enable crowd-sourced interactive evaluations, but they face validity issues by gamifying safety testing (e.g., with tasks like "break the model in one minute") and reducing safety to vulnerability against jailbreaking (Angelopoulos et al., 2024). Instead, we need more platforms that standardize the foundational elements of human subject studies with AI while maintaining scientific rigor, reducing technical overhead for researchers, and enabling broader participation in AI evaluation.
@@
How can interactive evaluations produce actionable findings that guide stakeholder decisions? What are the limitations of controlled studies in capturing broader impacts?
@@
Randomized control trials (RCTs) have proven valuable for policy decisions by providing rigorous evidence, particularly in public health and social policy (Hariton and Locascio, 2018). Similarly, interactive evaluations aim to produce actionable findings about how model behaviors influence human perception and behavior. These insights can help stakeholders make better decisions about model deployment, safety mechanisms, and interaction design. However, we currently lack adequate visibility into how different stakeholders — from AI labs to government bodies — actually use evaluation results in their decision-making processes. Bridging this gap requires closer collaboration between researchers and decision-makers to ensure evaluation designs align with practical needs (Hardy et al., 2024).
@@
Finally, we must be mindful of limitations; like RCTs, controlled evaluations of human-AI interaction may effectively measure individual-level effects (like individual manipulation or overreliance) while still missing broader systemic patterns. Thus, while measuring these individual impacts is crucial for understanding immediate safety concerns, they should be complemented with examinations of how AI systems reshape institutional structures, professional practices, and social arrangements — effects that emerge beyond individual interactions (Chater and Loewenstein, 2023).
@@ 6 Conclusion
As AI systems become increasingly integrated into our daily lives, moving beyond content risks to examine harms that emerge through sustained engagement is not just an academic exercise — it is essential for developing responsible technologies that prioritize human well-being. Several established industries, from medicine to automotives, have long recognized the need to study their technologies’ impact on users through extensive testing during development and after deployment (Wouters et al., 2020). As AI capabilities and applications continue to expand, we must similarly strengthen our investment in understanding these systems’ effects on human behavior and society at large.

### v7

@@ 4.2 Principle 2: identify the causal link between model behavior and human impact
These measurement targets can be assessed through both self-reported measures — including validated scales for attitudes, perceptions, and affective states — and behavioral measures that track concrete actions like response time, task accuracy, and engagement patterns (Damacharla et al. 2018; Coronado et al. 2022; Gordon et al. 2021). In Table 2, we present a non-exhaustive set of these metrics for assessing both human impact and human-AI performance.
@@
Psychological impact User affective states, or perceptions and attitudes toward the model or the interaction Usability metrics, user satisfaction surveys, psychometrically validated measures for a specific construct or harm (e.g., Decision Regret Scale (Brehaut et al. 2003))
@@
4.3 Principle 3: structure human participation to balance validity and practicality
@@
Chat logs from real-world interactions provide unparalleled ecological validity and scale. These naturally occurring conversations capture authentic user behaviors and emergent patterns that would be difficult to replicate in controlled settings (Tamkin et al. 2024). The massive volume of existing logs also enables analysis of rare events and edge cases that controlled studies might miss. Recent work has paired analyses of millions of chat logs of user conversations with ChatGPT with longitudinal surveys of those users to examine emotional well-being, showing a path forward for using chat logs while also assessing real user impact (Phang et al. 2025). However, this observational data is often proprietary, difficult to access, raises serious privacy concerns, and may limit our ability to test specific hypotheses or establish causal relationships (Reuel et al. 2024; Zhao et al. 2024; Ouyang et al. 2023).
@@
Existing research on user simulations has utilized LLMs to simulate believable human behaviors using a range of architectures from demographic-based models to simulations informed by qualitative interviews (Park et al. 2024, 2023). Recent work has used such simulations to evaluate safety risks, utilizing human experiments to validate the evaluations (Ibrahim et al. 2025; Zhou et al. 2024). Such simulations can be particularly valuable for exploring scenarios that would be ethically challenging to test with real users. Unlike real human participants, who naturally shape conversations in diverse ways, simulations also allow for controlled trajectories, making it easier to analyze how interactions evolve across different scenarios. However, their limitations in believability, accuracy, and diversity require further study, as recent work has shown that they misportray and flatten representations of marginalized identity groups (Wang, Morgenstern, and Dickerson 2025; Agnew et al. 2024).
@@ 5 Open challenges and ways forward for interactive evaluations
In this paper, we motivate the need to investigate human-AI interaction dynamics that current evaluations miss. However, implementing such evaluations at scale presents several concrete challenges. These include ethical questions about studying potentially harmful interactions, practical needs for research infrastructure, and methodological questions about producing actionable insights for stakeholders. While our design principles outline key considerations for scenarios, measurements, and participation strategies, advancing these methods requires addressing several open questions. Here, we identify some challenges and opportunities where concentrated research effort could significantly advance our ability to evaluate increasingly interactive AI systems.
@@
5.1 How can we ethically work with human participants on studying harms? When are user simulations appropriate replacements?
@@
Studying interaction harms poses inherent ethical challenges: we need to understand potentially harmful dynamics without exposing human participants to those same harms. This requires careful consideration of when and how to involve human participants. For participant protection, AI researchers should adopt existing practices from fields experienced in providing adequate participant training, thoughtful debriefing, and active feedback collection to mitigate experimental harms (McKee 2024). Simulated human interactions offer a promising direction for studying scenarios where minimizing harm is challenging, such as young people’s interactions with AI systems or emotional attachment to AI systems. Recent work suggests these simulations can believably represent human behavior (Park et al. 2023, 2024), but more work is needed to understand how to best capture diverse user behaviors and develop ethical guidelines for deploying user simulations in safety evaluations (Agnew et al. 2024; Anthis et al. 2025).
5.2 How can we improve researcher access to data for understanding interaction harms?
@@
Current public datasets of human-AI conversations, while valuable, fall short of meeting researchers’ needs. Large collections of chat logs like WildChat and ShareGPT offer millions of conversations but provide only a narrow window into AI usage, capturing self-selected interactions that may not represent typical user behaviors or patterns (Zhao et al. 2024; Ouyang et al. 2023). These datasets can be biased in terms of user demographics, use cases, and interaction styles, potentially leading to incomplete or incorrect conclusions about how people typically interact with AI systems. AI developers and service providers typically have access to the most comprehensive human-AI interaction data (Tamkin et al. 2024). Given their incentives to protect user privacy and proprietary information, developing privacy-preserving data-sharing frameworks is essential to enable researchers to access representative interaction data while maintaining user trust and consent.
5.3 What infrastructure do we need to facilitate interactive evaluations?
@@
Many interactive evaluations share common stages: recruiting human participants, collecting interaction data, and administering surveys and other tasks. Developing accessible protocols, guides, and standardized test suites — similar to those available for static evaluations — can support and facilitate an increase in interactive evaluations (UK AI Safety Institute 2024; METR 2023; Collins et al. 2023). Some existing platforms facilitate crowd-sourced interactive evaluations, but they may face validity concerns by gamifying safety testing (e.g., with tasks like ’break the model in one minute’) and focusing primarily on vulnerabilities such as jailbreaking (Angelopoulos et al. 2024). Instead, we need more platforms that standardize the foundational elements of human subject studies with AI while maintaining scientific rigor, reducing technical overhead for researchers, and enabling broader participation in AI evaluation.
Additionally, current platforms primarily support single-session studies, but understanding many interaction harms requires longitudinal evaluation capabilities with secure, privacy-preserving mechanisms for tracking behavioral changes over extended AI usage periods. This is a critical infrastructure gap that must be addressed to adequately assess the long-term influence of AI systems on human users.
@@
5.4 How can interactive evaluations produce actionable findings that guide stakeholder decisions? What are the limitations of controlled studies in capturing broader impacts?
@@
Randomized control trials (RCTs) have proven valuable for policy decisions by providing rigorous evidence, particularly in public health and social policy (Hariton and Locascio 2018). Similarly, interactive evaluations aim to produce actionable findings about how AI systems influence human perception and behavior. These insights can help stakeholders make better decisions about model deployment, safety mechanisms, and interaction design. However, we currently lack adequate visibility into how different stakeholders — from AI labs to government bodies — actually use evaluation results in their decision-making processes. Bridging this gap requires closer collaboration between researchers and decision-makers to ensure evaluation designs align with practical needs (Hardy et al. 2024).
@@
Finally, we must be mindful of limitations; like RCTs, controlled evaluations of human-AI interaction may effectively measure individual-level effects (like individual manipulation or overreliance) while still missing broader systemic patterns. Thus, while measuring these individual impacts is crucial for understanding immediate safety concerns, they should be complemented with examinations of how AI systems reshape institutional structures, professional practices, and social arrangements — effects that emerge beyond individual interactions (Chater and Loewenstein 2023).
@@ 6 Conclusion
As AI systems become increasingly integrated into our daily lives, examining harms that emerge through sustained engagement is essential for developing systems that prioritize human well-being. Our work makes three key contributions to this pursuit: first, we motivate the need to measure interaction harms and demonstrate why current evaluation paradigms systematically fail to capture them; second, we develop a structured framework of practical principles for designing interactive evaluations; and third, we identify specific implementation challenges and research directions that must be addressed for these methods to succeed at scale. Several established industries, from medicine to automotives, have long recognized the need to study their technologies’ impact on people through extensive testing during development and after deployment (Wouters, McKee, and Luyten 2020). As AI capabilities and applications continue to expand, we must similarly strengthen our investment in understanding these systems’ impact on human behavior and society at large. The methodological considerations outlined here provide a foundation for this shift in how we evaluate increasingly interactive AI systems.